The Term Titer in Medicine refers to the Concentration of Antibody Molecules or Virus Particles as determined by a Dilution Series. The Titer of our fX174 Bacteriophage is usually 1.0 x 1010 Plaque Forming Units per Milliliter (pfu/ml), indicating that there are at least 1.0 x 1010 Bacteriophage per Milliliter. Assuming that our fx 174 Bacteriophage Titer actually is 1.0 x 1010 Pfu/ml, and assuming that each Bacteriophage actually produces a Plaque, approximately how many Plaques should you expect to see on your 10-7 (17), 10-8 (8) and 10 9 (9) Phage Plates next Week? [1 Point] Focus IEK English (US) MacBook
Answer #1

Use the following formula to determine the titre of your viral stock:

Titer(pfu/ml)= No. of plaques/d*v

Where,

d= dilution

v= volume of diluted virus added to the plate

Here, in the question, we have to calculate no. of plaques.

Given,

Titer= 1.0 X 1010 PFU/ml

d=10-7

(i) For (10-7) =

1*1010pfu/ml= No. of plaques/10-7*1

=1*1010 *10-7= 103=1000 plaques

(ii) For 10-8

=1* 1010*10-8=102

=100 plaques

(iii) For 10-9

1*1010*10-9=101

=10 plaques